OpenStudy (anonymous):

Simplify by removing the factors of 1. z^2-16/4-z

OpenStudy (anonymous):

Im getting (z-4)(z+4)/(z-4) but I think I am missing a step

OpenStudy (anonymous):

You swapped the negative in your denominator. Which you can do (and need to in order to cancel properly) but you have to factor out a -1 in order to make it work. \[\frac{(z+4)(z-4)}{4-z} = \frac{(z+4){(z-4)}}{-1(z-4)} = -1(z+4) = -z-4 \text{ , } z \ne 4 \]

OpenStudy (anonymous):

thanks